OpenStudy (anonymous):

Factor completely 11s^2+392st+343t2

OpenStudy (anonymous):

i dont know but you could use the quadratic equation -b+or- sq.rt. of(b^2-4ac) all divided by 2a

OpenStudy (anonymous):

use a=343s, b=392s,c=11s t=[-b+- sq.rt(b^2-4ac)]/2a t=[-(392s)-+sqrt((-302s)^2 -4(343)(11s))]/2(343) t=[-(392s)-+sqrt((153664s^2-15092s^2)]/(686) t=[-392s-+sqrt(138572s^2)]/686 t=[-392s-+(14sqrt707)s]/686 t=[-392s-+372.25]/686 +... t=-285.64s/686 =-0.4163879s ans - ... t=-498.35788s/686=-0.726469s ans

OpenStudy (anonymous):

11s^2+392st+343t^2 =(t+0.726469s)(t-0.4163879s ) =(t-285.64s/686)(t-498.35788s/686) =((686t-285.64s)/686))((686t-498.35788s)/686) ans..